OpenStudy (anonymous):

Evaluate sec three pi over two to the nearest hundredth. The angle is given in radians

OpenStudy (anonymous):

do you know that \[\cos(\frac{3\pi}{2})\] is ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then take the reciprocal (flip it)

OpenStudy (anonymous):

secant is the reciprocal of cosine, so since \[\cos(\frac{3\pi}{2})=-\frac{1}{2}\] you have \[\sec(\frac{3\pi}{2})=-2\]

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

that was all it asked right?

OpenStudy (anonymous):

yeah but -2 isnt a choice lool